# Towards Practical Finite Sample Bounds for Motion Planning in TAMP

### v1

@@ A.4.2 Numerical Algorithm for Proposition 1 and Experimental Results
To verify the tightness of Prop. 1 and its numerical algorithm 5, we run an approximate $K$-nearest neighbors algorithm and compute the effective radius upper bound for varying numbers of neighbors, samples, and dimensions. We see that the bound often starts of conservative, and then progressively gets tighter as samples increase (with maximum number of samples of $n=10^{6}$. The rate at which the bound tightens in dependent on the dimension. The rate for $d=2,3$ is quite fast, while the does not fully connverge to tightly bound the right tail of the empirical distributions when $d=20$. Eventually, as $n\to\infty$, we expect the bound will become loose again.
Figure 5: The largest connection radius computed by the numerical bound by setting the probability of failure $\gamma=0.01$ is represent by the blue bars. Their right-to-left ordering is the same as the ordering of the histograms.

### v3

@@ A.4.2 Numerical Algorithm for Proposition 1 and Tightness Experiments
To verify the tightness of Prop. 1 and its numerical algorithm 5, we sample uniformly from a unit cube, run an approximate $K$-nearest neighbors algorithm, and compute the effective radius upper bound for varying numbers of neighbors, samples, and dimensions. The experiment is repeated for a variety of dimensions and nearest-neighbor counts. We see that the bound is overly conservative with small sample sets and then progressively tightens as samples increase (with a maximum number of samples of $n=10^{6}$). For $d=2,3$ the numerical bound tightens quickly but does not fully converge for $d=20$ within $10^{6}$ samples. Eventually, as $n\to\infty$, we expect the bound to become loose again because of the conservative logarithmic factor.
Figure 5: The largest connection radius computed by the numerical bound by setting the probability of failure $\gamma=0.01$ is represented by corresponding dotted vertical bars. Their right-to-left ordering is the same as the ordering of the histograms.
A.5 Code Implementation
The code that implements all the algorithms and experiments described in this paper is available in a public repository: https://github.com/robustrobotics/nonasymptotic-mp.